4. What is the probability that a that a worker at Random drives more that 15 miles to work if the distribution of all worker’s mileage is a Normal Distribution that ranges from 4 miles to 26 miles?
Answer by stanbon(75887)

What is the probability that a worker at Random drives more that 15 miles to work if the distribution of all worker’s mileage is a Normal Distribution that ranges from 4 miles to 26 miles?
----
range = 6 sigma
----
6 sigma = 26-4 = 22
---
sigma = 11/3
---
mean = 4+(22/2) = 15
-------------------------
P(x > 15) = 0.5000
========================
Cheers,
Stan H.